# Getting Bored of Cyberwar: Exploring the Role of Low-level Cybercrime Actors in the Russia-Ukraine Conflict

### v6

@@ 4. The Evidence from Web Defacements
There was a spike of 771 defacements by 5 defacers targeting Russia on 25 May. Of these, 764 were claimed by a single defacer compromising a server hosting 760 sites. This outlier appears to be unique; it was removed from the graph for better visualisation. The peak of 187 defacements hitting Ukraine on 1 February 2022 by 4 defacers did not have a single cause and did not lead to a sharp increase in defacers in the following days.
@@
For both Russia and Ukraine, Kruskal-Wallis tests suggest a statistically significant difference in the number of defacements and defacers per day through , , and , see Table 2 (Appendix §H). Post-hoc analysis indicates a significant difference between the pre-invasion () and one-month-post-invasion () periods. is also significant, but not , suggesting the situation returned back to the pre-war levels after the second era. The effect sizes are all between medium and large, ranging [0.06, 0.14].
@@
The Global Scale. The number of web defacement attacks targeting Russia and Ukraine appears to be trivial when set against the global scale. Among 274 963 analysed defacements, only 5 899 (2.15%) targeted the two countries (4 340 for Russia and 1 559 for Ukraine). The top 10 countries account for 69.85% of all defacements; sites hosted in the US by global-scale vendors (e.g., Cloudflare, Amazon) are excluded (see Appendix §C), but the US still consistently suffers the majority of web defacements. Since January 2022, the US accounts for 26.95% of defacement attacks, followed by India with 11.47% and Indonesia with 8.41%, while Russia and Ukraine only account for small proportions: 1.58% and 0.57%, respectively.
@@
There was a short-lived decline in defacement attacks worldwide on the invasion day (from around 1 400 to 1 000), while it peaked for Russia from nearly zero to 209 (14.48% of all defacements). This suggests a genuine change in the way defacers chose their targets, precipitated by the war. The US is consistently the largest target, but only accounts for 21.97% on that day. During the last two weeks of March 2022, the number of defacements significantly increased at the global scale, with many defacements targeting the US. However, much like the Russia-Ukraine scale, the effect lasted for only a few weeks. The unusual peaks against Brazil happened in late June (also for DDoS attacks, see Section §5), without a clear explanation.
@@
The patterns seen from the Kruskal-Wallis and post-hoc tests in the Russia-Ukraine scale does not apply for most top countries, see Table 2 (Appendix §H). ANOVA/Kruskal-Wallis tests on the defacement and defacer counts are not both significant for Brazil, Germany, India, and Singapore; no significant changes are seen between the pre-invasion and one-month-post-invasion eras for Canada and Turkey. Indonesia has a similar pattern of post-hoc tests, but one of the effect sizes is small. The only country following a close pattern is the US with medium effect sizes, yet Section §5 will show this did not hold for DDoS attacks on the US.
@@
The evidence above suggests a genuine increase of defacements against the two countries shortly after the invasion, standing out significantly from other top countries. Russia was the first to be hit at scale, followed by Ukraine a few days later. However, this effect was fairly short-lived for both countries, lasting for only a few weeks before returning to pre-war levels, presumably as defacers ran out of targets or just lost interest in carrying out attacks. The number of defacers involved was small, but for a while they turned from indiscriminate to more targeted attacks.
@@
We annotate motives based on 1 341 unique messages left on the defaced pages. We consider a political sentiment and mark it as supporting Russia/Ukraine if a support/objection is expressed e.g., ‘We stand with Ukraine!’. We mark messages consisting of defacers’ signature e.g., ‘Hacked by Hero’ without clear motives, or just greetings to peers as being for self-aggrandisement. Messages advertising hacking tools and services or asking for ransom are marked financially motivated e.g., ‘Contact me for shells’. We label messages expressing favourite mottos or moods as self-expression e.g., ‘Not much I want, hope my life will be better’, and exclude 1 278 messages (21.66%) containing empty or random messages.
@@ 5. The Evidence from DDoS Attacks
The Russia-Ukraine Scale. DDoS attacks lagged defacement by about a week, but occurred in higher volumes and lasted longer; most happened after 7AM. The number of both DDoS attacks and victims targeting Russia first increased on 2 March (6 days after the invasion) with 851 victims, 511 of them at around 6PM. The attacks peaked 4 days after with 1 137 victims. High activity levels continued through 23 March, with the biggest wave occurring at around 2PM on 8 March with 755 victims. Smaller waves continued regularly during the next few weeks. Regarding DDoS attacks hitting Ukraine, significant waves started around a week after Russia’s first big wave (some small spikes targeting Ukraine before Russia were insignificant) with the first notable spike on 10 March having 526 victims, then became prevalent during two weeks from 18 to 31 March: big waves were on 18 March at around 12PM, 1PM and 4PM with 257, 476, and 700 victims, respectively. Other big and medium waves lasted until the end of March, with the biggest peak on 31 March when 1 296 victims were hit. The increased volume only continued for about a month before declining sharply.
@@
The Global Scale. We again see concentrations in DDoS attacks, with the top 10 countries accounting for 70.49% of all victims. The US still dominates (24.68%), followed by Brazil (11.99%) and Bangladesh (8.10%). Ukraine took 1.57%, while Russia lies 8th at 3.61%. Our DDoS and defacements datasets show some correlations. Three of the top 10 countries for defacements are also in top 10 for DDoS targets: the US, Germany, and Brazil. As with defacements, the number of DDoS attacks rose globally during the last two weeks of March 2022. As with defacements, the DDoS attack counts rose globally during the last two weeks of March 2022 with large numbers targeting Bangladesh, but this is insignificant; see Table 3 (Appendix §H). The unusual peaks of both defacements and DDoS against Brazil in late June are notable; Brazil is often ranked among top cybercrime hubs worldwide (Lusthaus et al., 2020), yet we lack a convincing causality. A similar peak observed in the Russia-Ukraine scale can also be seen at the global scale following the invasion. As with defacements, DDoS attacks thrived on a global scale in March, yet they quickly returned to their previous levels after a few weeks with no lasting global changes and no significant evidence of further waves, whether targeting Russia and Ukraine.
The pattern seen from the Kruskal-Wallis and post-hoc tests in the Russia-Ukraine scale does not apply for most top countries, see Table 3 (Appendix §H). Kruskal-Wallis tests are not all significant for the US, Bangladesh, and Poland, despite the US accounting for the largest number of attacks and there was a visual increase for Bangladesh (as the tests compare mean ranks instead of means). No significant changes are seen between the pre-invasion and one-month-post-invasion periods for France and the UK. Brazil and Germany have a similar phenomenon of post-hoc tests in the Russia-Ukraine scale, yet one effect size is small. China is the only country following that phenomenon with large effect sizes; the main difference is that the changes in is not significant.
As with defacements, the evidence above suggests a genuine increase of DDoS attacks targeting Russia and Ukraine as the conflict began, standing out significantly from most top countries. Russia was still the first to be hit at scale, followed by Ukraine shortly after. The main difference with defacements is that DDoS attacks hitting Russia did not settle back to previous levels, but remained slightly higher. The outbreak of both defacement and DDoS attacks against Russia and Ukraine was significant and timely, but fairly short-lived: it returned to pre-war levels after just a few weeks.
@@ 6. Hacking Community Reactions
Hack Forums Discussions. There was an immediate surge of posts on Hack Forums mentioning the two countries after the invasion, from near zero to over 120 per day; see Figure 5. Kruskal-Wallis tests confirm the significance , with a large effect size ; pairwise post-hoc tests for and are both significant (), but not (). The number of posting users shows a similar story: Kruskal-Wallis test reports with a large effect size ; pairwise post-hoc tests for and are both significant (), but not (). This fits the evidence seen with defacement and DDoS attacks: both posting activity and users returned to the pre-war level after a few weeks, presumably as users lost interest and moved on to other discussion topics.
@@
Targets Promoted by the IT Army of Ukraine. Many announcements and targeted domains are posted in the first 2 weeks after the invasion, beginning on 26 February, peaking on 27 February with 40 announcements and 45 domains promoted (IP addresses were not regularly included until later), see Figure 6. Yet they quickly declined to consistently less than 10 per day after two weeks with some days (e.g., 24 and 26 April) having no posts. The number of subscribers also dropped from 300k to around 160k in October 2023.
@@
While the number of announcements dropped, the number of targets steadily increased, particularly in May and June 2022 with multiple-target posting. Activities were unstable at that time; targets got promoted less frequently and occasional days had no targets. Targets were mostly fresh in the first 2 weeks, but then a considerable proportion got re-promoted on multiple days e.g., all advertised IP addresses and most domains were re-posted during 4–6 May. Along with frequent zero-target days, this suggests the group might run out of new targets or get bored with finding them.
@@
Community reactions and engagement tell much the same story as with DDoS and defacement attacks (see Figure 7). While more targets were promoted in May and June, volunteers appeared to have largely lost interest, despite their intense activity in the first few weeks. The decline in reaction was consistent across all engagement types: views, emojis, forwards, and replies. Older announcements may have more time to accrue views as people scroll up the channel, but the emojis, forwards, and replies require user intent. We believe the figures reflect a genuine decline in engagement over time.
@@
We further looked at user engagement, with instructions about tools and guidance to carry out attacks provided by the group. The group first provided material to hit Russian payment systems on 9 March (two weeks after the invasion), attracting high levels of engagement: 240k views, 2.6k emojis, 1.2k forwards, and 421 replies from 197 users. The next was on 1 April: while the number of replies and forwards was similar to the first, other kinds roughly halved. From mid-May to late June, instructions were posted 4 more times, yet users were around 4 times less engaged than the first in March, indicating a loss of interest despite the operator’s extensive efforts.
@@
Target Selection. Targets were often themed, and sometimes patterned around particular weekdays. For instance, online news and propaganda, food delivery services and entertainment are often hit at weekends to maximise impact as people spend more time online. Themes were also occasionally set with re-promoted old targets, leading to wide variations in the number of new targets, particularly from May onwards (zero on some days), see Figure 6. Subscribers can suggest new targets, but the group owner posts most of them; during the last two months, they often re-promoted targets linked to old posts, which could be as simple as ‘we continue to work with yesterday’s targets’. This could suggest that the group organisers might have some difficulty in coming up with new targets.
@@
Categories vary, yet five dominate 80.21% of all targets, see Figure 8. ‘News, media and propaganda’, including TV broadcasting, has been consistently promoted since the war began, but only became the most common one in May when it overtook ‘IT solutions and services’. ‘Government and public services’, which includes military, state-owned websites, and public services for civilians such as parking and lighting (including governments imposed on occupied territories) has also been regularly targeted throughout, but they only grew rapidly towards the end of the period, making this the second most common category overall. ‘Banking and finance’ ranked 3rd, including banks, stock exchanges, electronic payment, accounting, credit services, trading, bidding, investment platforms and funding agencies. ‘IT solutions and services’ ranked 4th, including software solutions supporting governments, digital signature and information security services such as DDoS-Guard. This sector was actively promoted early on but was targeted far less thereafter. ‘Logistics’ ranks 5th, including airlines and aviation, travel, shipping, and food delivery. Other popular categories (by popularity) include markets and stores (e.g., job markets, real estate, e-commerce, drug stores), manufacturers and trading (e.g., military footwear, wood and roofing materials), education, insurance, telecommunications (e.g., Internet providers), businesses and state companies (e.g., energy and steel manufacturers), discussion forums, entertainment (e.g., cinemas), and non-governmental organisations.
@@
Crossover with Observed Attacks. The IT Army of Ukraine maintains a dashboard of targets’ status, claiming many are down due to their actions. To find whether the attacks involved reflected DDoS or defacement, we correlate our attacks records with promoted targets since the Telegram group started. We consider a defacement overlap when either its URL or IP address matches promoted targets, while for DDoS attacks, only IP addresses are used. There was very little overlap with website defacement attacks: among 3 845 targets promoted by the group, there are only 59 valid matches (1.53%), including 7 domain matches (0.18%) and 52 IP matches (1.35%). Notably, no overlaps occur on the day targets are promoted, suggesting that defacers chose their targets themselves independently (by scanning sites based on ccTLDs); these targets are largely unimportant and irrelevant to the conflict. For DDoS attacks, we observe 707 (30.86%) total overlaps among 2 291 promoted IP addresses, which is considerable. Unlike defacements, some are executed the same day they are promoted; we find many same-day overlaps in late March, early April and during May, peaking on 19 March 2022 with 22 victims overlapping. However, the crossover dropped quickly, becoming less frequent from late May while many new targets were still actively advertised. This suggests a loss of interest by volunteers in attacking targets promoted by the group.

### v7

@@ 4. The Evidence from Web Defacements
There was a spike of 771 web defacements by five defacers targeting Russia on 25 May. Of these, 764 were claimed by a single defacer compromising a server hosting 760 sites. This outlier appears to be unique; it was removed from the graph for better visualisation. The peak of 187 web defacements hitting Ukraine on 1 February 2022 by four defacers did not have a single cause and did not lead to a sharp increase in the number of defacers in the following days.
@@
For both Russia and Ukraine, Kruskal-Wallis tests suggest a statistically significant difference in the number of web defacements and defacers per day through , , and , see Table 2 (Appendix §H). Post-hoc analysis indicates a significant difference between the pre-invasion () and one-month-post-invasion () periods. is also significant, but not , suggesting the situation returned to the pre-war levels after the second era. The effect sizes are all between medium and large, ranging [0.06, 0.14].
@@
The Global Scale. The number of web defacement attacks targeting Russia and Ukraine appears to be trivial when set against the global scale. Among 274 963 analysed web defacements, only 5 899 (2.15%) targeted the two countries (4 340 for Russia and 1 559 for Ukraine). The top 10 countries account for 69.85% of all web defacements; sites hosted in the US by global-scale vendors (e.g., Cloudflare, Amazon) are excluded (see Appendix §C), but the US still consistently suffers the majority of web defacements. Since January 2022, the US accounts for 26.95% of defacement attacks, followed by India with 11.47% and Indonesia with 8.41%, while Russia and Ukraine only account for small proportions: 1.58% and 0.57%, respectively.
@@
There was a short-lived decline in defacement attacks worldwide on the invasion day (from around 1 400 to 1 000), while it peaked for Russia from nearly zero to 209 (14.48% of all defacements). This suggests a genuine change in the way defacers chose their targets, precipitated by the war. The US is consistently the largest target but only accounts for 21.97% on that day. During the last two weeks of March 2022, the number of defacements significantly increased at the global scale, with many defacements targeting the US. However, much like the Russia-Ukraine scale, the effect lasted for only a few weeks. The unusual peaks against Brazil happened in late June (also for DDoS attacks, see Section §5), without a clear explanation.
@@
The patterns seen from the Kruskal-Wallis and post-hoc tests in the Russia-Ukraine scale do not apply for most top countries, see Table 2 (Appendix §H). ANOVA/Kruskal-Wallis tests on the defacement and defacer counts are not both significant for Brazil, Germany, India, and Singapore; no significant changes are seen between the pre-invasion and one-month-post-invasion eras for Canada and Turkey. Indonesia has a similar pattern of post-hoc tests, but one of the effect sizes is small. The only country following a close pattern is the US with medium effect sizes, yet Section §5 will show this did not hold for DDoS attacks on the US.
@@
The evidence above suggests a genuine increase of website defacements against the two countries shortly after the invasion, standing out significantly from other top countries. Russia was the first to be hit at scale, followed by Ukraine a few days later. However, this effect was fairly short-lived for both countries, lasting for only a few weeks before returning to pre-war levels, presumably as defacers ran out of targets or just lost interest in carrying out attacks. The number of defacers involved was small, but for a while, they turned from indiscriminate to more targeted attacks.
@@
We annotate motives based on 1 341 unique messages left on the defaced pages. We consider a political sentiment and mark it as supporting Russia/Ukraine if a support/objection is expressed e.g., ‘We stand with Ukraine!’. We mark messages consisting of defacer signatures e.g., ‘Hacked by ABC’ without clear motives, or just greetings to peers as being for self-aggrandisement. Messages advertising hacking tools and services or asking for ransom are marked financially motivated e.g., ‘Contact me for shells’. We label messages expressing favourite mottos or moods as self-expression e.g., ‘Not much I want, hope my life will be better’, and exclude 1 278 messages (21.66%) containing empty or random messages.
@@ 5. The Evidence from DDoS Attacks
The Russia-Ukraine Scale. DDoS attacks lagged defacement by about a week, but occurred in higher volumes and lasted longer; most happened after 7 AM. The number of both DDoS attacks and victims targeting Russia first increased on 2 March (six days after the invasion) with 851 victims, 511 of them at around 6 PM. The attacks peaked four days later with 1 137 victims. High activity levels continued through 23 March, with the biggest wave occurring at around 2 PM on 8 March with 755 victims. Smaller waves continued regularly during the next few weeks. Regarding DDoS attacks hitting Ukraine, significant waves started around a week after Russia’s first big wave (some small spikes targeting Ukraine before Russia were insignificant) with the first notable spike on 10 March having 526 victims, then became prevalent during two weeks from 18 to 31 March: big waves were on 18 March at around 12 PM, 1 PM and 4 PM with 257, 476, and 700 victims, respectively. Other big and medium waves lasted until the end of March, with the biggest peak on 31 March when 1 296 victims were hit. The increased volume only continued for about a month before declining sharply.
@@
The Global Scale. We again see concentrations in DDoS attacks, with the top 10 countries accounting for 70.49% of all victims. The US still dominates (24.68%), followed by Brazil (11.99%) and Bangladesh (8.10%). Ukraine took 1.57%, while Russia lies 8th at 3.61%. Our DDoS and defacement datasets show some correlations. Three of the top 10 countries for defacements are also in the top 10 for DDoS targets: the US, Germany, and Brazil. As with defacements, DDoS attacks rose globally during the last two weeks of March 2022 with large numbers targeting Bangladesh, but this effect is insignificant; see Table 3 (Appendix §H). The unusual peaks of both defacements and DDoS against Brazil in late June are notable; Brazil is often ranked among the top cybercrime hubs worldwide (Lusthaus et al., 2020), yet we lack a convincing causality. A similar peak observed on the Russia-Ukraine scale can also be seen at the global scale following the invasion. As with defacements, DDoS attacks thrived on a global scale in March, yet quickly returned to the previous levels after a few weeks with no lasting global changes and no significant evidence of further waves, whether targeting Russia or Ukraine.
The patterns seen from the Kruskal-Wallis and post-hoc tests in the Russia-Ukraine scale do not apply for most top countries, see Table 3 (Appendix §H). Kruskal-Wallis tests are not all significant for the US, Bangladesh, and Poland, despite the US accounting for the largest number of attacks and there was a visual increase for Bangladesh (as the tests compare mean ranks instead of means). No significant changes are seen between the pre-invasion and one-month-post-invasion periods for France and the UK. Brazil and Germany have a similar phenomenon of post-hoc tests in the Russia-Ukraine scale, yet one effect size is small. China is the only country following that phenomenon with large effect sizes; the main difference is that the changes in are not significant.
As with defacements, the evidence above suggests a genuine increase in DDoS attacks targeting Russia and Ukraine as the conflict began, standing out significantly from most top countries. Russia was still the first to be hit at scale, followed by Ukraine shortly after. The outbreak of both defacement and DDoS attacks on Russia and Ukraine was significant and timely, but fairly short-lived. While defacements returned to previous levels after a few weeks, DDoS attacks on Russia remained marginally higher than pre-war levels.
@@ 6. Hacking Community Reactions
Hack Forums Discussions. There was an immediate surge of posts on Hack Forums mentioning the two countries after the invasion, from near zero to over 120 per day; see Figure 5. Kruskal-Wallis tests confirm the significance , with a large effect size ; pairwise post-hoc tests for and are both significant (), but not (). The number of posting users shows a similar story: Kruskal-Wallis test reports with a large effect size ; pairwise post-hoc tests for and are both significant (), but not (). This fits the evidence seen with defacement and DDoS attacks: both posting activity and users returned to the pre-war level after a few weeks, presumably as users quickly lost interest and moved on to other discussion topics.
@@
Targets Promoted by the IT Army of Ukraine. Many announcements and targeted domains were posted in the first two weeks after the invasion, beginning on 26 February, peaking on 27 February with 40 announcements and 45 domains promoted (IP addresses were not regularly included until later), see Figure 6. Yet they quickly declined to consistently less than 10 per day after two weeks with some days (e.g., 24 and 26 April) having no posts. The number of subscribers also dropped from 300k to around 160k in October 2023.
@@
While the number of announcements dropped, the number of targets steadily increased, particularly in May and June 2022 with multiple-target posting. Activities were unstable at that time; targets got promoted less frequently and occasional days had no targets. Targets were mostly fresh in the first two weeks, but then a considerable proportion got re-promoted on multiple days e.g., all advertised IP addresses and most domains were re-posted during 4–6 May. Along with frequent zero-target days, this suggests the group might run out of new targets or get bored with finding them.
@@
Community reactions and engagement tell much the same story as with DDoS and defacement attacks (see Figure 7). While more targets were promoted in May and June, volunteers appeared to have largely lost interest, despite their intense activity in the first few weeks. The decline in reaction was consistent across all engagement types: views, emojis, forwards, and replies. Older announcements may have more time to accrue views as people scroll up the channel, but the emojis, forwards, and replies require user intent. We believe the figures reflect a genuine decline in user engagement over time.
@@
The group first provided instructions about tools and guidance to carry on attacks against Russian payment systems on 9 March (two weeks after the invasion), attracting high levels of engagement: 240k views, 2.6k emojis, 1.2k forwards, and 421 replies from 197 users. The next was on 1 April: while the number of replies and forwards was similar to the first, other kinds roughly halved. From mid-May to late June, instructions were posted four more times, yet users were around four times less engaged than the first in March, indicating a loss of interest despite the operator’s intensive efforts.
@@
Target Selection. Targets were often themed, patterning around particular weekdays e.g., online news and propaganda, food delivery, and entertainment are often attacked at weekends to maximise impact as people spend more time online. Themes were also occasionally set with re-promoted old targets, leading to wide variations in the number of new targets, particularly from May onwards (zero on some days, see Figure 6). Subscribers can suggest new targets, but group organisers post most of them. During the last two months, they often re-promoted targets linked to old posts, which could be as simple as ‘we continue to work with yesterday’s targets’, suggesting that they might have some difficulties in hunting new targets.
@@
Categories vary, yet five dominate 80.21% of all targets, see Figure 8. ‘News, media and propaganda’, including TV broadcasting, has been consistently promoted since the war began but only became the most common one in May when it overtook ‘IT solutions and services’. ‘Government and public services’, which includes military, state-owned websites, and public services for civilians such as parking and lighting (including governments imposed on occupied territories) has also been regularly targeted throughout, but they only grew rapidly towards the end of the period, making this the second most common category overall. ‘Banking and finance’ ranked 3rd, including banks, stock exchanges, electronic payment, accounting, credit services, credit card providers, trading, bidding, investment platforms, funding agencies, and the microfinance industry. ‘IT solutions and services’ ranked 4th, including software solutions supporting governments, digital signature and information security services such as DDoS-Guard. This sector was actively promoted early on but was targeted far less thereafter. ‘Logistics’ ranks 5th, including airlines and aviation, travel, shipping, and food delivery. Other categories by popularity include (online) marketplaces and stores (e.g., job markets, real estate, e-commerce, drug stores, app stores), manufacturers and trading (e.g., military footwear, shoes, wood and roofing materials), education, insurance, telecommunications (e.g., Internet providers), businesses and state companies (e.g., energy and steel manufacturers), online forums, entertainment (e.g., cinemas), and non-governmental organisations.
@@
Crossover with Observed Attacks. The IT Army of Ukraine maintains a dashboard of targets’ status, claiming many are down due to their actions. To find whether the attacks involved reflected DDoS or defacement, we correlate our attack records with promoted targets since the Telegram group started. We consider a defacement overlap when either its URL or IP address matches promoted targets, while for DDoS attacks, only IP addresses are used. There was very little overlap with website defacement attacks: among 3 845 targets promoted by the group, there are only 59 valid matches (1.53%), including 7 domain matches (0.18%) and 52 IP matches (1.35%). Notably, no overlaps occur on the day targets are promoted, suggesting that defacers chose their targets themselves independently (by scanning sites based on ccTLDs); these targets are largely unimportant and irrelevant to the conflict. For DDoS attacks, we observe 707 (30.86%) total overlaps among 2 291 promoted IP addresses, which is considerable. Unlike defacements, some are executed the same day they are promoted; we find many same-day overlaps in late March, early April, and during May, peaking on 19 March 2022 with 22 victims overlapping. However, the crossover dropped quickly, becoming less frequent from late May while many new targets were still actively advertised. This suggests a loss of interest by volunteers in attacking targets promoted by the group.